### v1

@@ Abstract
Recently, Álvarez-Rebollar, Cravioto-Lagos, Marín, Solé-Pi, and Urrutia (Graphs and Combinatorics, 2024) constructed a set of points for which the longest perfect matching is noncrossing. They raised several challenging questions in this direction. In particular, they asked whether the longest spanning path, on any finite set of points in the plane, must have a pair of crossing edges. They also conjectured that the longest spanning cycle must have a pair of crossing edges.
@@ 1 Introduction
Traversing points in the plane by a polygonal path or cycle possessing a desired property has a rich background. For instance, the celebrated travelling salesperson problem asks for a polygonal path or cycle with minimum total edge length [9, 25, 27]. In recent years, there has been increased interest in paths and cycles with properties such as being noncrossing [2, 17], minimizing the longest edge length [6, 12, 24], maximizing the shortest edge length [7], minimizing the total or largest turning angle [1, 14, 19, 22], and minimizing the number of turns [13, 18, 28] to name a few. The longest cycle—the MaxTSP—is NP-hard in Euclidean spaces of dimension $\geq 3$, but the complexity of the planar MaxTSP is unknown [10, 11, 21]. [^1] It is interesting that the MaxTSP under the $L_{1}$-norm for points in the plane can be solved in linear time [11, 21], in contrast to the fact that MinTSP in this case is NP-hard [23]. Paths and cycles that have combinations of these properties have also attracted attention. For example, simultaneously being noncrossing and having maximum total edge length [3, 20] is difficult to satisfy: to achieve a larger length we typically introduce more crossings.
@@
Edge crossings in geometric graphs are usually undesirable as they have the potential of creating unwanted situations such as collisions in motion planning and inconsistency in VLSI layout. They are also undesirable in the context of graph drawing and network visualization as they make drawings more difficult to read and use. Short geometric structures such as shortest perfect matchings, shortest spanning trees, shortest spanning paths, and shortest spanning cycles are inherently noncrossing. This property, however, does not necessarily hold if the structure is not shortest. For long structures such as longest perfect matchings, longest spanning trees, longest spanning paths, and longest spanning cycles—the other end of the spectrum—it seems natural to expect many crossings. Counting crossings in geometric graphs and finding geometric structures with a minimum or maximum number of crossings are active research areas in discrete geometry. The study of this type of problem attracted more attention after the work of Aronov et al. [8] in 1994, who showed that any set of $n$ points in the plane in general position admits a crossing family (a set of pairwise intersecting segments) of size $\Omega(\sqrt{n})$. They also conjectured that the true lower bound is linear in $n$. The current best lower bound, $n^{1-o(1)}$, was established by Pach et al. [26] in 2019.
@@
The noncrossing property of shortest structures is mainly ensured by the triangle inequality. The triangle inequality, as noted by Alon et al. [3], also implies that the longest structures often have crossings because a structure usually gets longer by creating more crossings. Alon et al. [3] studied the problem of finding longest noncrossing structures (such as matchings, paths, or trees). Some of their initial results have been improved and extended by Dumitrescu and Tóth [20] (for matchings, paths, and cycles), by Biniaz et al. [15] and by Cabello et al. [16] (for trees). Along this direction, one might wonder whether a longest structure (defined on an arbitrarily large point set) is necessarily crossing. This was explicitly asked by Álvarez-Rebollar et al. [4]. Among other interesting results, they presented arbitrarily large planar point sets for which the longest perfect matching is noncrossing. They asked the following question and proposed the following conjecture:
Question 1 (Álvarez-Rebollar et al. [4]).
@@
The “sufficiently large” condition in the question and conjecture makes sense, as otherwise one can take any 3 points in general position, or any $4$ points that are not in a convex position—for such point sets, all spanning paths and cycles are noncrossing.
In the other direction, one might wonder about maximizing the number of crossings in cycles. Here, we would like to highlight another result of Álvarez-Rebollar et al. [4, 5]. Let $C(n)$ be the largest number such that any set of $n$ points in the plane admits a spanning cycle with at least $C(n)$ pairs of crossing edges. Álvarez-Rebollar et al. [4, 5] established the following lower and upper bounds: $n^{2}/12-O(n)<C(n)<5n^{2}/18-O(n)$. In other words, any set of $n$ points in the plane admits a spanning cycle with at least $n^{2}/12-O(n)$ crossings, and there is a family of point sets that does not admit any cycle with more than $5n^{2}/18-O(n)$ crossings.
@@ 1.1 Our contributions
In this paper, we provide negative answers to both Question 1 and Conjecture 1. For any integer $n\geq 1$ we present a set of $n$ points in the plane for which the longest spanning path is unique and noncrossing. Similarly, for any integer $n\geq 4$, we present a set of $n$ points in the plane for which the longest spanning cycle is unique and noncrossing. To build such point sets, we use the following framework: First, we choose a set $P$ of points on the $x$-axis for which the longest structure may not be unique. Then, we assign new $y$-coordinates to points in $P$ to obtain a new point set $P^{\prime}$ for which the longest structure corresponds to one in $P$ and is also unique and noncrossing. In Section 6, we present some structural properties of longest paths and cycles, which may be of independent interest.
@@ 1.2 Preliminaries
Consider two line segments, each connecting a pair of points in $P$. If the interiors of the segments intersect, then we say that they cross; this configuration is called a crossing. A path or a cycle is called noncrossing if its edges do not cross each other. We denote the undirected edge between two points $p$ and $q$ by $pq$, the directed edge from $p$ towards $q$ by $(p,q)$, and the Euclidean distance between $p$ and $q$ by $|pq|$. The length of a geometric graph $G$ is the sum of the lengths of its edges, and we denote it by $|G|$.
@@ 2 Longest Paths and Cycles on the Real Line
In this section we characterize longest paths and cycles in dimension one. These observations play a pivotal role in our constructions in the plane (Sections 3 and 4). We say that an edge $e$ intersects a point $p$ if the intersection of $e$ and $p$ is not empty (the intersection could be an endpoint of $e$). For a sorted set of $12k{+}$ 1 numbers, the median is the number with rank $k{+}1$, and for a sorted set of $2k$ numbers, the median is the mean of the two numbers with ranks $k$ and $k{+}1$.
@@
Let $P$ be a set with an even number of points in $\mathbb{R}$, i.e., in dimension one. The endpoints of any longest spanning path on $P$ lie on different sides of the median of $P$.
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ and assume w.l.o.g. that 0 is the median of $P$ (in particular, $0\notin P$). Let $H$ be a longest spanning path on $P$. Orient the edges of $H$ to make it a directed path. Let $p_{s}$ and $p_{e}$ be the starting and ending points of $H$, respectively. For the sake of contradiction, assume that $p_{s}$ and $p_{e}$ have the same sign, which we may assume, due to symmetry, to be positive. Thus $p_{s},p_{e}>0$. Then, the sum of degrees of vertices in $H$ to the left of the origin is 2 more than the sum of degrees of vertices to the right. Therefore, $H$ must have a directed edge $(p_{a},p_{b})$ where $p_{a},p_{b}<0$. If $p_{b}<p_{a}$, then by replacing $(p_{a},p_{b})$ with the undirected edge $p_{s}p_{b}$ we obtain a longer undirected path; and if $p_{b}>p_{a}$ by replacing $(p_{a},p_{b})$ with $p_{e}p_{a}$ we obtain a longer undirected path. Both cases lead to a contradiction. ∎
@@
Let $P$ be a set with an even number of points in $\mathbb{R}$, i.e., in dimension one. Let $H$ be a spanning path on $P$. Then $H$ is a longest spanning path if and only if
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ so that $p_{i}<p_{j}$ for all $i<j\in\{1,\ldots,n\}$, and assume w.l.o.g. that 0 is the median of $P$. Note that $0\notin P$ since $n$ is even. First, we prove by contradiction that if $H$ is a longest spanning path, then (i) and (ii) hold.
Suppose that (i) does not hold. Orient the edges of $H$ to make it a directed path. Let $(p_{a},p_{b})$ be an edge of $H$ that does not intersect the median. Due to symmetry, assume that $p_{a},p_{b}<0$. By Lemma 1, the endpoints of $H$ lie on different sides of the median. This implies that both sides have the same sum of vertex degrees. Thus $H$ must have an edge $(p_{c},p_{d})$ such that $c,d>0$. By replacing these edges with $p_{a}p_{c}$ and $p_{b}p_{d}$ we obtain an (undirected) spanning path that is longer than $H$ because $|p_{a}-p_{c}|+|p_{b}-p_{d}|>|p_{a}-p_{b}|+|p_{c}-p_{d}|$. This contradicts $H$ being a longest path.
Now suppose that (ii) does not hold: without loss of generality $p_{n/2}$ is not an endpoint of $H$. (The case for $p_{n/2+1}$ can be handled symmetrically). Then $H$ has an endpoint $p_{a}$ with $a<n/2$. Orient the edges of $H$ so that the path is directed from $p_{a}$ towards the other endpoint. Let $(p_{n/2},p_{b})$ be the outgoing edge from $p_{n/2}$. By part (i), we have $p_{b}\geq 0$. By removing $(p_{n/2},p_{b})$ we obtain two paths, and $p_{b}$ is an endpoint on one of those paths. Next, join the paths with a new edge $(p_{a},p_{b})$. Thus we obtain an (undirected) spanning path that is longer than $H$ because $|p_{a}-p_{b}|>|p_{n/2}-p_{b}|$. This contradicts $H$ being longest.
Finally, we prove that any spanning path $H$ that satisfies (i) and (ii) is longest, using a direct proof. Consider a longest spanning path $L$ on $P$. By the sufficiency proof, (i) and (ii) hold for $L$. This implies that the positive interval $[p_{n/2},p_{n/2+1}]$ is contained in each of the $n{-}1$ edges, hence it contributes to the length of $L$ with multiplicity $n-1$. Similarly, for any $i\in\{2,\ldots,n/2\}$ the positive interval $[p_{n/2+i-1},p_{n/2+i}]$ contributes to the length of $L$ by multiplicity $n-2i$. A similar argument holds for negative intervals. See Figure 1. On the other hand, any spanning path (including $H$) that satisfies (i) and (ii) receives the exact same multiplicities from the corresponding intervals. Therefore $H$ and $L$ have the same length, and hence $H$ is also a longest path. ∎
Figure 1: Illustration of a longest path for a point set on a line, for the case where the number of points, $n$, is even. Numbers below intervals $[p_{n/2+i},p_{n/2+i+1}]$ represent the multiplicity of the contribution of the corresponding intervals to the length of the longest path.
A statement similar to that of Lemma 2 can be proved for paths with an odd number of points (in this case one endpoint is the median itself and the other endpoint is the closest point to the median). However, we will not use this in our construction.
@@
Let $P$ be a finite set in $\mathbb{R}$, i.e., in dimension one.
@@
    A spanning cycle on $P$ is longest iff each of its edges intersects the median of $P$.
@@
    If $P$ contains an odd number of points, then for any longest spanning cycle the two edges incident to the median lie on opposite sides of it.
@@
[…]
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ so that $p_{i}<p_{j}$ for all $i<j\in\{1,\ldots,n\}$, and assume w.l.o.g. that 0 is the median of $P$. Note that $0\notin P$ if $n$ is even, and $p_{\lceil n/2\rceil}=0$ if $n$ is odd.

### v2

@@ Abstract
Recently, Álvarez-Rebollar, Cravioto-Lagos, Marín, Solé-Pi, and Urrutia (Graphs and Combinatorics, 2024) constructed a set of points for which the longest perfect matching is noncrossing. They raised several challenging questions in this direction. In particular, they asked whether the longest spanning path, on every finite set of points in the plane, must have a pair of crossing edges. They also conjectured that the longest spanning cycle must have a pair of crossing edges.
@@ 1 Introduction
Traversing points in the plane by a polygonal path or cycle possessing a desired property has a rich background. For instance, the celebrated travelling salesperson problem asks for a polygonal path or cycle with minimum total edge length [9, 27, 29]. In recent years, there has been increased interest in paths and cycles with properties such as being noncrossing [2, 18], minimizing the longest edge length [6, 12, 25], maximizing the shortest edge length [7], minimizing the total or largest turning angle [1, 13, 20, 23], and minimizing the number of turns [14, 19, 30] to name a few. Finding a longest cycle—the MaxTSP problem—is NP-hard in Euclidean spaces of dimension three or higher, but the complexity of the MaxTSP problem in the plane is unknown [10, 11, 22]. [^1] It is interesting that the MaxTSP under the $L_{1}$-norm for points in the plane can be solved in linear time [11, 22], in contrast to the fact that MinTSP in this case is NP-hard [24]. Paths and cycles that have combinations of these properties have also attracted attention. For example, maximizing the total edge length subject to the constraint that the edges are noncrossing [3, 21] is difficult: to achieve a larger length we typically introduce more crossings.
@@
Edge crossings in geometric graphs are usually undesirable as they have the potential of creating unwanted situations such as collisions in motion planning and inconsistency in VLSI layout. They are also undesirable in the context of graph drawing and network visualization as they make drawings more difficult to read and use. Short est geometric structures such as shortest perfect matchings, shortest spanning trees, shortest spanning paths, and shortest spanning cycles are inherently noncrossing. This property, however, does not necessarily hold if the structure is not shortest. For long structures such as longest perfect matchings, longest spanning trees, longest spanning paths, and longest spanning cycles—the other end of the spectrum—it seems natural to expect many crossings. Counting crossings in geometric graphs and finding geometric structures with a minimum or maximum number of crossings are active research areas in discrete geometry. This line of research is also related to geometric Ramsey-type problems [8, 26]. For example, it is known that every set of $n$ points in the plane in general position admits a crossing family (a matching with pairwise crossing edges) of size $n^{1-o(1)}$ [28], but it remains open whether this bound can be improved to $\Omega(n)$ [16].
@@
The noncrossing property of shortest structures is mainly ensured by the triangle inequality. The triangle inequality also implies that the longest structures often have crossings because a structure usually gets longer by creating more crossings  [3]. There are approximation algorithms for the problem of finding longest noncrossing structures (such as matchings, paths, cycles, and trees) [3, 15, 17, 21]. Along this direction, one might wonder whether a longest structure (defined on an arbitrarily large point set) is necessarily crossing. This was explicitly asked by Alvarez-Rebollar, Cravioto-Lagos, Marín, Solé-Pi, and Urrutia [4]. Among other interesting results, they presented arbitrarily large planar point sets for which the longest perfect matching is noncrossing. They asked the following question and proposed the following conjecture:
Question 1 ( [4]).
@@
The “sufficiently large” condition in the question and conjecture cannot be dropped, as otherwise one can take any 3 points in general position, or any 4 points that are not in a convex position—for such point sets, all spanning paths and cycles are noncrossing.
In the other direction, one might wonder about maximizing the number of crossings in cycles. Let $C(n)$ be the largest number such that every set of $n$ points in the plane admits a spanning cycle with at least $C(n)$ pairs of crossing edges. The following lower and upper bounds are known  [4, 5]: $\frac{1}{12}\,n^{2}-O(n)<C(n)<\frac{5}{18}\,n^{2}-O(n)$. In other words, every set of $n$ points in the plane admits a spanning cycle with at least $\frac{1}{12}\,n^{2}-O(n)$ crossings, and there are arbitrarily large point sets that do not admit any cycle with more than $\frac{5}{18}\,n^{2}-O(n)$ crossings.
@@ 1.1 Our Contributions
In this paper, we provide a negative answer to Question 1 and refute Conjecture 1. For every integer $n\geq 1$, we present a set of $n$ points in the plane for which the longest spanning path is unique and noncrossing. Similarly, for every integer $n\geq 3$, we present a set of $n$ points in the plane for which the longest spanning cycle is unique and noncrossing. To build such point sets, we use the following framework: First, we choose a set $P$ of points on the $x$-axis for which the longest structure may not be unique. Then, we assign new $y$-coordinates to points in $P$ to obtain a new point set $P^{\prime}$ for which the longest structure corresponds to one in $P$ and is both unique and noncrossing.
The paper is organized as follows. In Section 2 we prove structural properties of longest paths and cycles on the real line. We employ these properties and construct noncrossing longest paths in Section 3, noncrossing longest cycles in Section 4, and noncrossing longest matchings in Section 5. In Section 6, we present some structural properties of longest paths and cycles in the plane, which may be of independent interest.
@@ 1.2 Preliminaries
Consider two line segments, each connecting a pair of points in $P$. If the relative interiors of the segments intersect, then we say that they cross; this configuration is called a crossing. A path or a cycle is called noncrossing if no two edges cross. We say that an edge $e$ intersects the $y$-axis if the intersection of $e$ and the $y$-axis is not empty (the intersection could be an endpoint of $e$). We denote the undirected edge between two points $p$ and $q$ by $pq$, the directed edge from $p$ towards $q$ by $(p,q)$, and the Euclidean distance between $p$ and $q$ by $|pq|$. The length of an edge with endpoints $p$ and $q$ is the Euclidean distance between $p$ and $q$. The length of a geometric graph $G$ is the sum of the lengths of its edges, and we denote it by $|G|$. In general, if $w:E\to\mathbb{R}$ is a weight function on the edges of a graph $G=(V,E)$, then the weight of a subgraph $H$ of $G$ is the sum of the weights of its edges: $w(H)=\sum_{e\in E(H)}w(e)$.
The following lemma and its corollary are simple but play an important role in our constructions for noncrossing longest paths and cycles.
Lemma 1.
Let $G=(V,E)$ be a graph with two edge weight functions $w_{1}:E\to\mathbb{R}$ and $w_{2}:E\to\mathbb{R}$ such that for every $e\in E$, we have $|w_{1}(e)-w_{2}(e)|\leq 2\,\alpha$. If $\mathcal{A}$ and $\mathcal{B}$ are two families of subgraphs of $G$ with $m$ edges such that $\beta:=\min\{w_{1}(A)-w_{1}(B):A\in\mathcal{A},B\in\mathcal{B}\}>0$, then $\min\{w_{2}(A)-w_{2}(B):A\in\mathcal{A},B\in\mathcal{B}\}\geq\beta-4m\alpha$.
Proof.
For every $A\in\mathcal{A}$ and $B\in\mathcal{B}$, we have
$w_{2}(A)-w_{2}(B)$ $=\sum_{e\in E(A)}w_{2}(e)-\sum_{f\in E(B)}w_{2}(f)$
$\geq\sum_{e\in E(A)}\big(w_{1}(e)-|w_{2}(e)-w_{1}(e)|\big)-\sum_{f\in E(B)}\big(w_{1}(f)+|w_{2}(f)-w_{1}(f)|\big)$
$\geq\sum_{e\in E(A)}w_{1}(e)-\sum_{f\in E(B)}w_{1}(f)-4m\alpha$
$=w_{1}(A)-w_{1}(B)-4m\alpha.$
Consequently, we have $\min\{w_{2}(A)-w_{2}(B):A\in\mathcal{A},B\in\mathcal{B}\}\geq\beta-4m\alpha$. ∎
Corollary 1.
Let $P$ be a set of $n$ points in the plane and $\mathcal{S}$ a family of geometric graphs on $P$ with at most $m$ edges (for example, spanning cycles with $m=n$ or spanning paths with m=n1m=n-1). Let $\mathcal{S}=\mathcal{A}\cup\mathcal{B}$ be a partition such that $\mathcal{A}\neq\emptyset$ and $\beta:=\min\{|A|-|B|:A\in\mathcal{A},B\in\mathcal{B}\}>0$. Assume that every point in $P$ is perturbed by a distance of at most $\alpha$, $0\leq\alpha<\frac{\beta}{4m}$, and let $\mathcal{S}^{\prime}$ be the set of perturbed graphs corresponding to $\mathcal{S}$. Then every longest graph in $\mathcal{S}^{\prime}$ corresponds to a graph in $\mathcal{A}$.
Proof.
We use Lemma 1 with the following setup: Let $G=(P,E)$ be the complete graph on $P$. For every edge $pq\in E$, let $w_{1}(pq)=|pq|$ and $w_{2}(pq)=|p^{\prime}q^{\prime}|$, where $p$ is perturbed to $p^{\prime}$ and $q$ to $q^{\prime}$. If every point is perturbed by a distance of at most $\alpha$, then for every edge $pq\in E$ the triangle inequality yields $w_{2}(pq)=|p^{\prime}q^{\prime}|\leq|pp^{\prime}|+|pq|+|qq^{\prime}|\leq w_{1}(pq)+2\alpha$ and $w_{2}(pq)=|p^{\prime}q^{\prime}|\geq|pq|-|pp^{\prime}|-|qq^{\prime}|\geq w_{1}(pq)-2\alpha$, hence $|w_{2}(pq)-w_{1}(pq)|\leq 2\alpha$.
Now consider a longest graph $G^{\prime}\in\mathcal{S}^{\prime}$, corresponding to a graph $G\in\mathcal{S}$. We have $|G^{\prime}|=w_{2}(G)$ and $|G|=w_{1}(G)$. If $G\notin\mathcal{A}$, then for every graph $A\in\mathcal{A}$, Lemma 1 gives $w_{2}(A)-w_{2}(G)\geq\beta-4\alpha m>0$, which means that $G^{\prime}$ is not a longest graph in $\mathcal{S}^{\prime}$: a contradiction. We conclude that $G\in\mathcal{A}$, as required. ∎
@@ 2 Longest Paths and Cycles on the Real Line
In this section we characterize the longest paths and cycles in dimension one. These observations play a pivotal role in our constructions in the plane (Sections 3 and 4). We say that an edge $e$ intersects a point $p$ if the intersection of $e$ and $p$ is not empty (the intersection could be an endpoint of $e$). For a sorted set of 2k+12k{+}1 numbers, the median is the number at position k+1k{+}1, and for a sorted set of $2k$ numbers, the median is the mean of the two numbers at positions $k$ and k+1k{+}1.
@@
Let $P$ be a set of $n\geq 2$ points in $\mathbb{R}$, and let $H$ be a longest spanning path on $P$. Then.
  • $(i)$
    the two endpoints cannot lie on strictly the same side of the median of $P$, and
  • $(ii)$
    every edge of $H$ intersects the median of $P$.
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ and assume without loss of generality that 0 is the median of $P$ . Note that $0\in P$ if $n$ is odd and $0\notin P$ if $n$ is even. We prove both statements (i) and (ii) by contradiction.
Suppose that (i) does not hold. Orient the edges of $H$ to make it a directed path. Let $p_{s}$ and $p_{e}$ be the starting and ending points of $H$, respectively. For the sake of contradiction, assume that $p_{s}$ and $p_{e}$ have the same sign, which we may assume, due to symmetry, to be positive, that is, $p_{s},p_{e}>0$. Then, the sum of degrees of vertices in $H$ to the left of the origin is 2 more than the sum of degrees of vertices to the right. Therefore, $H$ must have a directed edge $(p_{a},p_{b})$ where $p_{a},p_{b}\leq 0$. If $p_{b}<p_{a}$, then by replacing the undirected edge $p_{a}p_{b}$ with the edge $p_{b}p_{s}$ we obtain a longer undirected path; and if $p_{b}>p_{a}$ by replacing $p_{a}p_{b}$ with $p_{e}p_{a}$ we obtain a longer undirected path. Both cases lead to a contradiction.
Suppose that (ii) does not hold. Let $(p_{a},p_{b})$ be an edge of $H$ that does not intersect the median. Due to symmetry, we may assume that $p_{a},p_{b}<0$. Part (i) implies that the sum of degrees on negative and positive vertices differ by at most one (and must be equal when $n$ is even). Thus $H$ must have an edge $(p_{c},p_{d})$ such that $p_{c},p_{d}\geq 0$. By replacing these edges with $p_{a}p_{c}$ and $p_{b}p_{d}$ we obtain an (undirected) spanning path that is longer than $H$ because $|p_{a}-p_{c}|+|p_{b}-p_{d}|>|p_{a}-p_{b}|+|p_{c}-p_{d}|$. This contradicts $H$ being a longest path. ∎
We can characterize longest spanning paths on an even number of points.
@@
Let $P$ be a set with an even number of points in $\mathbb{R}$. Let $H$ be a spanning path on $P$. Then $H$ is a longest spanning path if and only if
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ so that $p_{i}<p_{j}$ for all $\in\{1,\ldots,n\}, i<ji<j$, and assume without loss of generality that 0 is the median of $P$. Note that $0\notin P$ since $n$ is even. First, we prove that if $H$ is a longest spanning path on $P$, then (i) and (ii) hold. Property (i) follows from Lemma 2(ii). We prove (ii) by contradiction.
Suppose that (ii) does not hold: without loss of generality $p_{n/2}$ is not an endpoint of $H$. (The case for $p_{n/2+1}$ can be handled symmetrically). Then $H$ has an endpoint $p_{a}$ with $a<n/2$. Orient the edges of $H$ so that the path is directed from $p_{a}$ towards the other endpoint. Let $(p_{n/2},p_{b})$ be the outgoing edge from $p_{n/2}$. By property (i), we have $p_{b}>0$. By removing the undirected edge $p_{n/2}p_{b}$ we obtain two undirected paths, and $p_{b}$ is an endpoint on one of those paths. Next, join the paths with a new edge $p_{a}p_{b}$. Thus we obtain an spanning path that is longer than $H$ because $|p_{a}-p_{b}|>|p_{n/2}-p_{b}|$. This contradicts $H$ being longest.
Finally, we prove that every spanning path $H$ that satisfies (i) and (ii) is longest, using a direct proof. Consider a longest spanning path $L$ on $P$. By the necessity proof, (i) and (ii) hold for $L$. This implies that the interval $p_{n/2},p_{n/2+1}]$ is contained in each of the n1n{-}1 edges, hence it contributes to the length of $L$ with multiplicity n1n-1. Similarly, for every $i\in\{1,2,\ldots,n/2-1\}$ the intervals $p_{n/2-i},p_{n/2-i+1}]$ and $p_{n/2+i},p_{n/2+i+1}]$ each contribute to the length of $L$ with multiplicity n+22in+2-2i. See Figure 1. On the other hand, every spanning path (including $H$) that satisfies (i) and (ii) receives the exact same multiplicities from the corresponding intervals. Therefore $H$ and $L$ have the same length, and hence $H$ is also a longest path. ∎
Figure 1: Illustration of a longest path for a point set on a line, for the case where the number of points, $n$, is even. Numbers below intervals $p_{n/2+i},p_{n/2+i+1}]$ represent the multiplicity of the contribution of the corresponding intervals to the length of the longest path.
For an odd number of points, a similar argument yields essentially the same characterization as Lemma 3.
@@
Let $P$ be a set with an odd number of points in $\mathbb{R}$. Let $H$ be a spanning path on $P$. Then $H$ is a longest spanning path if and only if
@@
    every edge of $H$ intersects the median of $P$, and
@@
    one endpoint of $H$ is the median and the other endpoint is a closest point to the median.
Proof.
Let $n=2k+1$ and $P=\{p_{1},\ldots,p_{n}\}$ so that $p_{i}<p_{j}$ for all $\in\{1,\ldots,n\}, i<ji<j$. Note that the median of $P$ is $p_{k+1}$, and assume without loss of generality that $p_{k+1}=0$. We prove that if $H$ is a longest spanning path on $P$, then (i) and (ii) hold. Property (i) follows from Lemma 2(ii). We prove (ii) by contradiction.
First, suppose that the median $p_{k+1}=0$ is not an endpoint of $H$. By Lemma 2(i), the sum of degrees on (strictly) positive and negative vertices are the same. Consequently, the median is incident to a vertex on each side. Without loss of generality, $H$ contain the edges $(p_{a},0)$ and $(0,p_{b})$ where $p_{a}<0<p_{b}$. Replace the undirected edges $p_{a}0$ and $0p_{b}$ with $p_{a}p_{b}$ and $p_{e}0$, where $p_{e}$ is the endpoint of $H$. Since $|p_{a}|+|p_{b}|=|p_{b}-p_{a}|<|p_{b}-p_{a}|+|p_{e}|$, we obtain a spanning path longer than $H$, contradicting the maximality of $H$.
Second, suppose that the endpoints of $H$ are $p_{a}$ and $p_{k+1}=0$, but there is another point in $P$ between $p_{a}$ and 0. Assume without loss of generality that $p_{a}<p_{b}<0$. Orient the edges of $H$ so that the path is directed from $p_{a}$ towards the other endpoint 0. Let $(p_{b},p_{c})$ be the outgoing edge from $p_{b}$. By property (i), we have $p_{c}\geq 0$. Replace the undirected edge $p_{b}p_{c}$ with $p_{a}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{c}}$, to obtain a new spanning path on $P$, which is longer than $H$ since $|p_{b}-p_{c}|<|p_{a}-p_{c}|$, contradicting the maximality of $H$.
Finally, suppose that the endpoints of $H$ are $p_{a}$ and $p_{k+1}=0$, there is no point in $P$ between $p_{a}$ and 0, but the closest point to 0 is $p_{b}$ with $p_{b}\neq p_{a}$. Without loss of generality, assume $p_{a}<0<p_{b}$. Orient the edges of $H$ so that the path is directed from $p_{a}$ towards the other endpoint 0. Let $(p_{b},p_{c})$ be the outgoing edge from $p_{b}$, and $(p_{d},0)$ the last edge of $H$. By property (i), we have $p_{c}\leq 0$ and $p_{d}>0$. If $p_{b}p_{c}=p_{d}0$, then replace this edge $p_{b}p_{c}=p_{d}0=p_{b}0$ with $p_{a}0$ to obtain a new spanning path on $P$, which is longer since $|p_{b}|<|p_{a}|$. Otherwise $p_{b}p_{c}\neq p_{d}0$, which implies under our assumptions that $p_{c}<p_{a}<0<p_{b}<p_{d}$. In this case, replace the undirected edges $p_{b}p_{c}$ and $p_{d}0$ with $p_{a}p_{d}$ and $p_{c}0$, to obtain a new spanning path on $P$, which is longer than $H$ because $|p_{b}-p_{c}|+|p_{d}|=|p_{a}-p_{c}|+|p_{a}|+2\,|p_{b}|+|p_{d}-p_{b}|<|p_{a}-p_{c}|+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\,|p_{a}|}+|p_{b}|+|p_{d}-p_{b}|=|p_{a}-p_{d}|+|p_{c}|$, contradicting the maximality of $H$.
The sufficiency of conditions (i) and (ii) can be proved similarly to the proof of Lemma 3: comparing the contributions of the intervals $p_{i},p_{i+1}]$, $i=1,\ldots,n$, in a longest spanning path $L$ on $P$, and a spanning path satisfying (i) and (ii). If $p_{k}$ and $p_{k+1}$ are the endpoints of $H$, then the intervals $p_{k},p_{k+1}]$ and $p_{k+1},p_{k+2}]$ contribute to the length of $H$ with multiplicity n2n-2 and n1n-1, respectively. In the special case that the median has two closest points in $P$ (that is, $|p_{k}|=|p_{k+2}|$), then $p_{k+2}$ may be an endpoint of $H$, and then these intervals contribute n1n-1 and n2n-2, respectively. For every $i\in\{1,\ldots,k-1\}$, however, the intervals $p_{k-i},p_{k-i+1}]$ and $p_{k+i+1},p_{k+i+2}]$ each contribute with multiplicity 2(ki)2(k-i). ∎
We can also characterize longest spanning cycles on $n$ points on a line.
Lemma 5.
Let $P$ be a finite set in $\mathbb{R}$.
  • $(i)$
    A spanning cycle on $P$ is longest if and only if each of its edges intersects the median of $P$.
  • $(ii)$
    If $P$ contains an odd number of points, then for every longest spanning cycle the two edges incident to the median lie on opposite sides of it.
@@
    Assume that $P$ contains n=2k+1n=2k{+}1 points and there is an interval $I$ of length $h>0$ between the leftmost k+1k{+}1 and the rightmost $k$ points. Then in every longest spanning cycle, n1=2kn{-}1=2k edges contain the interval $I$; and if a spanning cycle has fewer than $2k$ edges that contain $I$, then it is at least $2h$ shorter than a longest cycle.
@@
Let $P=\{p_{1},\ldots,p_{n}\}$ so that $p_{i}<p_{j}$ for all $i<j\in\{1,\ldots,n\}$, and assume without loss of generality that 0 is the median of $P$. Note that $0\notin P$ if $n$ is even, and $p_{\lceil n/2\rceil}=0$ if $n$ is odd.